# Pronunciation Editing for Finnish Speech using Phonetic Posteriorgrams

### v2

@@ Abstract
Synthesizing second-language (L2) speech is potentially highly valued for L2 language learning experience and feedback. However, due to the lack of L2 speech synthesis datasets, it is difficult to synthesize L2 speech for low-resourced languages. In this paper, we provide a practical solution for editing native speech to approximate L2 speech and present PPG2Speech, a diffusion-based multispeaker Phonetic-Posteriorgrams-to-Speech model that is capable of editing a single phoneme without text alignment. We use Matcha-TTS’s flow-matching decoder as the backbone, transforming Phonetic Posteriorgrams (PPGs) to mel-spectrograms conditioned on external speaker embeddings and pitch. PPG2Speech strengthens the Matcha-TTS’s flow-matching decoder with Classifier-free Guidance (CFG) and Sway Sampling. We also propose a new task-specific objective evaluation metric, the Phonetic Aligned Consistency (PAC), between the edited PPGs and the PPGs extracted from the synthetic speech for editing effects. We validate the effectiveness of our method on Finnish, a low-resourced, nearly phonetic language, using approximately 60 hours of data. We conduct objective and subjective evaluations of our approach to compare its naturalness, speaker similarity, and editing effectiveness with TTS-based editing. Our source code is published at https://github.com/aalto-speech/PPG2Speech.
@@ 1 Introduction
Proficiency in second languages (L2) enhances personal communication skills and cultural understanding, and it is also an essential criterion for evaluating immigrants’ integration into the local society. Pronunciation, as a challenging yet essential part of language learning, benefits from pronunciation feedback from a teacher [1]. Recent studies demonstrate that speech synthesis technology effectively supports L2 pronunciation training by providing individual pronunciation feedback [2], and providing synthesized L2 pronunciation to the learner improves learning experience and results [3].
@@
Speech editing is a technique that allows insertion, deletion, and replacement of certain speech content. Current speech editing models [5, 6, 7, 8] adopt a mask-and-infill approach, which consists of masking the edited region, filling this region with new content, and ensuring prosody consistency by imposing auxiliary loss on the edge of the edited region. These methods are text-based, operate at the word-level, and depend on external aligners to provide alignment information to determine the edited region. Early speech editing methods [5, 6] are trained with reconstruction loss on the mel-spectrogram, making it less ideal in synthesis quality. Recent editing methods [7, 8] are diffusion-based, improving the synthesized sample quality.
@@
Diffusion-based Speech Synthesis using Flow Matching [9] has become a popular research topic in recent years [10, 11, 12]. As a class of Non-Autoregressive (NAR) models, flow-matching-based Speech Synthesis models benefit from parallel computation of all temporal steps and can effectively balance the synthesis quality and the inference speed by adjusting the number of diffusion sampling steps. However, NAR models require modeling the alignment between input text and output speech. Voicebox [10] and NaturalSpeech 3 [11] rely on an external phoneme-level aligner that expands the input phoneme sequence to the length of the synthesized speech based on the duration of each phoneme, before inputting the phoneme sequence to the model. Matcha-TTS [12] uses monotonic alignment search to find the alignment between phoneme representations and synthesized speech, and trains a duration predictor to predict the duration of each phoneme. All these methods [10, 11, 12] impose hard alignments between input text and synthesized speech, which might hurt the naturalness of the synthesized speech [13].
@@
To gain more control over the pronunciation information, we use Phonetic Posteriorgrams (PPGs) as the input, instead of the phoneme sequence. PPGs are time-varying categorical distributions over phonemes [14]. They have been widely used for different speech processing tasks [15, 16, 17, 18, 19, 20] in the past years. As a time-varying distribution over phoneme categories, they contain rich pronunciation information that makes them ideal for pronunciation assessment and mispronunciation detection [15], and spoken term detection [16]. PPGs also disentangle well with the timbre and excitation of speech, making them ideal as content representations for speech and singing voice conversion [17, 18, 19]. Furthermore, they are interpretable representations for humans and can be potentially used for neural speech editing [20].
@@
In this paper, we present PPG2Speech, a diffusion-based multispeaker PPG-to-speech model using Matcha-TTS’s flow-matching decoder [12], enhanced with classifier-free guidance (CFG) [21] and sway sampling [22]. Our proposed approach operates without textual input, providing a practical solution for editing native speech to approximate common mispronunciations made by L2 learners. Unlike the previous approaches in Speech Editing and TTS, which require explicit hard alignment for model training and determining the edited region, our model uses PPGs, which provide soft, explicit alignment information that retains the naturalness of synthetic speech. We validate our method with both objective and subjective evaluations to assess naturalness, speaker similarity, and editing effectiveness in comparison with TTS-based editing techniques. Objective metrics include a novel Phonetic Aligned Consistency (PAC) to assess editing effectiveness.
@@ 2.1 Flow matching for generative models
Flow Matching (FM) is a recent generative modeling technique designed to transform a simple prior distribution, such as a standard Gaussian, into a complex data distribution through deterministic flows defined by ordinary differential equations (ODEs) [9]. Given an unknown data distribution $q(x)$, FM constructs a probability density path $p_{t}(x)$, with $t\in[0,1]$, smoothly transitioning from a known prior distribution $p_{1}(x)=\mathcal{N}(x;0,I)$, a standard Gaussion distribution, to a final distribution $p_{0}(x)$ approximating the data $q(x)$.
@@
$L_{\text{OT-CFM}}(\theta)=\mathbb{E}_{t,q(x_{1}),p_{0}(x_{0})}\left[\\ ||v_{t}(\phi_{t}^{\text{OT}}(x);\theta)-(x_{0}-x_{1})\\ ||^{2}\right].$ (4)
@@ 2.1.1 Classifier-free guidance
Classifier-free guidance is a technique developed to enhance sample fidelity in conditional diffusion models without relying on an external classifier [21]. The key idea is to jointly train a conditional and an unconditional diffusion model using a shared neural network. During training, the model is conditioned on the input conditions $c$ with some probability $P_{\text{con}}$, and with a complementary probability $P_{\text{uncon}}=1-P_{\text{con}}$, it is trained unconditionally by providing a null condition. This allows the network to learn both conditional and unconditional score estimates $v_{t}(\phi_{t}(x_{t}),c)$ and $v_{t}(\phi_{t}(x_{t}))$, respectively.
@@ 2.1.2 Sway Sampling
Typically, $n$ timesteps are sampled uniformly for $t\in[0,1]$ during the diffusion sampling process, where $n$ is the total number of iterations in the process. Recent research [22] demonstrates the use of Sway Sampling improves the quality of the sample. The Sway Sampling is formulated as:
@@ 3.1 PPG2Speech
We show our PPG2Speech model’s architecture in Figure 1. Our model generates mel-spectrograms from PPGs with conditional pitch sequences and speaker embeddings. The model consists of a PPG encoder and a flow-matching decoder. The PPG encoder consists of a prenet, a stack of Transformer/Conformer [23] layers, with an upsampling layer in between. The prenet maps the PPG channels to the Transformer’s channel dimensions. The upsampling layer upsamples the latent PPG representations to the time resolution of the target mel-spectrogram based on the nearest neighbor method. The encoder uses Relative Positional Encoding [24] to provide positional information to the self-attention.
@@
For the flow-matching decoder, we extend the 1-D U-Net based decoder in Matcha-TTS [12] with Classifier-free Guidance (CFG). Furthermore, the decoder is conditioned on external speaker embeddings and pitch embedding sequences to generalize to unseen speakers and synthesize more natural speech. The original speaker embedding table in Matcha-TTS is discarded. During training, we uniformly sample the timestep $t\in[0,1]$ to predict the gradient vector field. During inference, we use Sway Sampling with $s=-1$ to transform the speech from the noise using smaller steps in the initial steps to improve the quality of the synthesized speech. The synthesized mel-spectrogram is converted to waveform using a pretrained HiFi-GAN vocoder [25].
@@ 4.1 Data
We conduct our experiments on the combination of two Finnish datasets, Perso Synteesi [26] and Finsyn [27]. We show the details of these two datasets in Table 1. Perso Synteesi [26] is a read-speech dataset collected from 33 male native speakers and 33 female native speakers by recording speakers’ voices when reading prompts. Prompts are the same for all speakers. We only have the data of 20 male speakers and 30 female speakers. Finsyn [27] is another Finnish speech synthesis dataset with two female speakers and various speaking styles. We follow the suggestion from [27] and filter out those utterances with a spontaneous speaker style to ensure stable synthesis results.
@@
We split our data into Train, Validation, Test, and Test_unseen sets. For the Perso Synteesi [26], we use the original partition of the dataset, and exclude speakers $19$, $20$, $19m$, $20m$ from the training and validation set and use their original validation and test set as Test_unseen for testing the model’s generalization on unseen speakers. For the Finsyn [27] dataset, we split the data into training, validation, and testing sets with ratio $0.9:0.05:0.05$. No utterance from Finsyn is split to Test_unseen because Finsyn only has 2 speakers. The resulting splits are shown in Table 2.
We extract the PPGs from the Kaldi HMM-TDNN-medium model reported in [28], which contains 29 Finnish phonemes, and 3 special symbols, SPN, SIL, and eps. We use the pretrained SimAMResNet34111https://github.com/wenet-e2e/wespeaker/blob/master/docs/pretrained.md model from WeSpeaker toolkit [29] to extract the speaker embeddings, and use the pretrained Pitch Estimation Neural Network (PENN) [30] to extract pitch and periodicity from the audio. We standardize the log pitch per utterance by subtracting its mean and dividing its variance to mitigate timbre leakage, and then quantize pitch into 256 bins. The mel-spectrograms are extracted from the audio using HiFi-GAN [25] configurations.
@@
Table 3: The objective results. SECS is the Speaker Encoder Cosine Similarity, the Character Error Rate (CER) is used to evaluate intelligibility. Furthermore, Mel-Cepstral Distortion (MCD) and pitch mean absolute error (Pitch MAE) are used to evaluate the model’s capability on unseen speakers. The Phonetic Aligned Consistency (PAC) is the objective evaluation for editing effects.
@@ 4.2 Model details
We first use a 3-layer convolutional prenet with each layer having a kernel size 3 to map the 32-dim PPGs to 128-dim latent. We use the 2-layer Conformer [23] before the upsampling layer with $4$ attention heads, 512 feed-forward dimension, and a convolution kernel size of 9. After the upsampling, we use a 2-layer Transformer with 4 attention heads and 512 feed-forward dimensions. We find this configuration yields the best outcome.
Our flow-matching decoder follows the same configuration as in [12]. The external speaker embedding is of 256 dimensions. We use an embedding table of dimension 16 to represent the quantized pitch and concatenate the log periodicity with the pitch embedding as the soft voiced/unvoiced (V/UV) flag. External speaker embedding and pitch embedding with the V/UV flag are concatenated together as the conditional input to the decoder. During training, there is a $10\%$ probability for each batch that the PPG latent and the conditional input will be dropped to train the unconditioned model. During inference, we set the guidance strength $w=3$, and the sway coefficient is set to $-1$.
We train both the Matcha-TTS and our PPG2Speech model from scratch using the training set and evaluate on the test set and the unseen test set. Both models are trained with a vanilla flow-matching decoder in Matcha-TTS and a flow-matching decoder strengthened by CFG, to show the CFG’s effect on the model’s performance. All models are trained with a batch size of 32 for 500k updates on 1 Nvidia A100 GPU. The maximum learning rate is 1e-4, and we use a cosine annealing learning rate scheduler with 150k steps linear warmup to adjust the learning rate during training.
@@ 4.3.1 Objective Evaluation
The results of the objective evaluation are shown in Table 3. We use Speaker Encoder Cosine Similarity (SECS) to measure the similarity between the groundtruth speaker embedding and the speaker embedding extracted from the synthesized speech using the same SimAMResNet34 model in Section 4.1. We use a Wav2Vec2-large model [^2] https://huggingface.co/GetmanY1/wav2vec2-large-fi-150k-finetuned that is pretrained on 158k hours of unlabeled Finnish speech and fine-tuned on 4600 hours of Finnish speech to evaluate the intelligibility of the synthesized speech and use Character Error Rate (CER) as the objective metric. Furthermore, we use Mel-Cepstral Distortion (MCD) [31] and pitch mean absolute error (Pitch MAE) to evaluate the model’s capability of handling unseen speakers and how well the model follows the given pitch, respectively.
@@
    CFG improves the model’s generalization to unseen speakers.
@@
    Estimate the PPG using the same HMM-TDNN-medium model as Section 4.1 and use Montreal-Forced-Aligner [32] to find the corresponding edited region. Calculate the PAC in Section 3.2 between the edited PPG and the estimated PPG in the edited region.

### v3

@@ Abstract
Synthesizing second-language (L2) speech is potentially highly valued for L2 language learning experience and feedback. However, due to the lack of L2 speech synthesis datasets, it is difficult to synthesize L2 speech for low-resourced languages. In this paper, we provide a practical solution for editing native speech to approximate L2 speech and present PPG2Speech, a diffusion-based multispeaker Phonetic-Posteriorgrams-to-Speech model that is capable of editing a single phoneme without text alignment. We use Matcha-TTS's flow-matching decoder as the backbone, transforming Phonetic Posteriorgrams (PPGs) to mel-spectrograms conditioned on external speaker embeddings and pitch. PPG2Speech strengthens the Matcha-TTS's flow-matching decoder with Classifier-free Guidance (CFG) and Sway Sampling. We also propose a new task-specific objective evaluation metric, the Phonetic Aligned Consistency (PAC), between the edited PPGs and the PPGs extracted from the synthetic speech for editing effects. We validate the effectiveness of our method on Finnish, a low-resourced, nearly phonetic language, using approximately 60 hours of data. We conduct objective and subjective evaluations of our approach to compare its naturalness, speaker similarity, and editing effectiveness with TTS-based editing. Our source code is published at https://github.com/aalto-speech/PPG2Speech.
@@ 1 Introduction
Proficiency in second languages (L2) enhances personal communication skills and cultural understanding, and it is also an essential criterion for evaluating immigrants' integration into the local society. Pronunciation, as a challenging yet essential part of language learning, benefits from pronunciation feedback from a teacher [Derwing2015]. Recent studies demonstrate that speech synthesis technology effectively supports L2 pronunciation training by providing individual pronunciation feedback [Liakin19052017], and providing synthesized L2 pronunciation to the learner improves learning experience and results [cardoso2018learning].
@@
Speech editing is a technique that allows insertion, deletion, and replacement of certain speech content. Current speech editing models [tan2021editspeech, wang2022campnet, jiang2023fluentspeech, liu2024fluenteditor] adopt a mask-and-infill approach, which consists of masking the edited region, filling this region with new content, and ensuring prosody consistency by imposing auxiliary loss on the edge of the edited region. These methods are text-based, operate at the word-level, and depend on external aligners to provide alignment information to determine the edited region. Early speech editing methods [tan2021editspeech, wang2022campnet] are trained with reconstruction loss on the mel-spectrogram, making it less ideal in synthesis quality. Recent editing methods [jiang2023fluentspeech, liu2024fluenteditor] are diffusion-based, improving the synthesized sample quality.
@@
Diffusion-based Speech Synthesis using Flow Matching [lipman2023flow] has become a popular research topic in recent years [le2023voicebox, ju2024naturalspeech, matchatts]. As a class of Non-Autoregressive (NAR) models, flow-matching-based Speech Synthesis models benefit from parallel computation of all temporal steps and can effectively balance the synthesis quality and the inference speed by adjusting the number of diffusion sampling steps. However, NAR models require modeling the alignment between input text and output speech. Voicebox [le2023voicebox] and NaturalSpeech 3 [ju2024naturalspeech] rely on an external phoneme-level aligner that expands the input phoneme sequence to the length of the synthesized speech based on the duration of each phoneme, before inputting the phoneme sequence to the model. Matcha-TTS [matchatts] uses monotonic alignment search to find the alignment between phoneme representations and synthesized speech, and trains a duration predictor to predict the duration of each phoneme. All these methods [le2023voicebox, ju2024naturalspeech, matchatts] impose hard alignments between input text and synthesized speech, which might hurt the naturalness of the synthesized speech [eskimez2024e2].
@@
To gain more control over the pronunciation information, we use Phonetic Posteriorgrams (PPGs) as the input, instead of the phoneme sequence. PPGs are time-varying categorical distributions over phonemes [ppgs]. They have been widely used for different speech processing tasks [6639269, 5372889, 7552917, miyoshi2017voice, 9414137, morrison2024fine] in the past years. As a time-varying distribution over phoneme categories, they contain rich pronunciation information that makes them ideal for pronunciation assessment and mispronunciation detection [6639269], and spoken term detection [5372889]. PPGs also disentangle well with the timbre and excitation of speech, making them ideal as content representations for speech and singing voice conversion [7552917, miyoshi2017voice, 9414137]. Furthermore, they are interpretable representations for humans and can be potentially used for neural speech editing [morrison2024fine].
@@
In this paper, we present PPG2Speech, a diffusion-based multispeaker PPG-to-speech model using Matcha-TTS's flow-matching decoder [matchatts], enhanced with classifier-free guidance (CFG) [ho2021classifierfree] and sway sampling [chen-etal-2024-f5tts]. Our proposed approach operates without textual input, providing a practical solution for editing native speech to approximate common mispronunciations made by L2 learners. Unlike the previous approaches in Speech Editing and TTS, which require explicit hard alignment for model training and determining the edited region, our model uses PPGs, which provide soft, explicit alignment information that retains the naturalness of synthetic speech. We validate our method with both objective and subjective evaluations to assess naturalness, speaker similarity, and editing effectiveness in comparison with TTS-based editing techniques. Objective metrics include a novel Phonetic Aligned Consistency (PAC) to assess editing effectiveness.
@@ 2.1 Flow matching for generative models
Flow Matching (FM) is a recent generative modeling technique designed to transform a simple prior distribution, such as a standard Gaussian, into a complex data distribution through deterministic flows defined by ordinary differential equations (ODEs) [lipman2023flow]. Given an unknown data distribution $q(x)$, FM constructs a probability density path $p_{t}(x)$, with $t\in[0,1]$, smoothly transitioning from a known prior distribution $p_{1}(x)=\mathcal{N}(x;0,I)$, a standard Gaussion distribution, to a final distribution $p_{0}(x)$ approximating the data $q(x)$.
@@
[…]
@@ 2.1.1 Classifier-free guidance
Classifier-free guidance is a technique developed to enhance sample fidelity in conditional diffusion models without relying on an external classifier [ho2021classifierfree]. The key idea is to jointly train a conditional and an unconditional diffusion model using a shared neural network. During training, the model is conditioned on the input conditions $c$ with some probability $P_{\text{con}}$, and with a complementary probability $P_{\text{uncon}}=1-P_{\text{con}}$, it is trained unconditionally by providing a null condition. This allows the network to learn both conditional and unconditional score estimates $v_{t}(\phi_{t}(x_{t}),c)$ and $v_{t}(\phi_{t}(x_{t}))$, respectively.
@@ 2.1.2 Sway Sampling
Typically, $n$ timesteps are sampled uniformly for $t\in[0,1]$ during the diffusion sampling process, where $n$ is the total number of iterations in the process. Recent research [chen-etal-2024-f5tts] demonstrates the use of Sway Sampling improves the quality of the sample. The Sway Sampling is formulated as:
@@ 3.1 PPG2Speech
We show our PPG2Speech model's architecture in Figure 1. Our model generates mel-spectrograms from PPGs with conditional pitch sequences and speaker embeddings. The model consists of a PPG encoder and a flow-matching decoder. The PPG encoder consists of a prenet, a stack of Transformer/Conformer [gulati2020conformer] layers, with an upsampling layer in between. The prenet maps the PPG channels to the Transformer's channel dimensions. The upsampling layer upsamples the latent PPG representations to the time resolution of the target mel-spectrogram based on the nearest neighbor method. The encoder uses Relative Positional Encoding [dai2019transformer] to provide positional information to the self-attention.
@@
For the flow-matching decoder, we extend the 1-D U-Net based decoder in Matcha-TTS [matchatts] with Classifier-free Guidance (CFG). Furthermore, the decoder is conditioned on external speaker embeddings and pitch embedding sequences to generalize to unseen speakers and synthesize more natural speech. The original speaker embedding table in Matcha-TTS is discarded. During training, we uniformly sample the timestep $t\in[0,1]$ to predict the gradient vector field. During inference, we use Sway Sampling with s=1s=-1 to transform the speech from the noise using smaller steps in the initial steps to improve the quality of the synthesized speech. The synthesized mel-spectrogram is converted to waveform using a pretrained HiFi-GAN vocoder [kong2020hifi].
@@ 4.1 Data
We conduct our experiments on the combination of two Finnish datasets, Perso Synteesi [perso] and Finsyn [finsyn]. We show the details of these two datasets in Table 1. Perso Synteesi [perso] is a read-speech dataset collected from 33 male native speakers and 33 female native speakers by recording speakers' voices when reading prompts. Prompts are the same for all speakers. We only have the data of 20 male speakers and 30 female speakers. Finsyn [finsyn] is another Finnish speech synthesis dataset with two female speakers and various speaking styles. We follow the suggestion from [finsyn] and filter out those utterances with a spontaneous speaker style to ensure stable synthesis results.
@@
We split our data into Train, Validation, Test, and Test_unseen sets. For the Perso Synteesi [perso], we use the original partition of the dataset, and exclude speakers $19$, 2020, $19m$, $20m$ from the training and validation set and use their original validation and test set as Test_unseen for testing the model's generalization on unseen speakers. For the Finsyn [finsyn] dataset, we split the data into training, validation, and testing sets with ratio $0.9:0.05:0.05$. No utterance from Finsyn is split to Test_unseen because Finsyn only has 2 speakers. The resulting splits are shown in Table 2.
We extract the PPGs from the Kaldi HMM-TDNN-medium model reported in [kaldi_model], which contains 29 Finnish phonemes, and 3 special symbols, SPN, SIL, and eps. We use the pretrained SimAMResNet34111https://github.com/wenet-e2e/wespeaker/blob/master/docs/pretrained.md model from WeSpeaker toolkit [wang2023wespeaker] to extract the speaker embeddings, and use the pretrained Pitch Estimation Neural Network (PENN) [morrison2023cross] to extract pitch and periodicity from the audio. We standardize the log pitch per utterance by subtracting its mean and dividing its variance to mitigate timbre leakage, and then quantize pitch into 256 bins. The mel-spectrograms are extracted from the audio using HiFi-GAN [kong2020hifi] configurations.
@@
Table 3: The objective results. SECS is the Speaker Encoder Cosine Similarity, the Character Error Rate (CER) is used to evaluate intelligibility. Furthermore, Mel-Cepstral Distortion (MCD) and pitch mean absolute error (Pitch MAE) are used to evaluate the model's capability on unseen speakers. The Phonetic Aligned Consistency (PAC) is the objective evaluation for editing effects.
@@ 4.2 Model details
We first use a 3-layer convolutional prenet with each layer having a kernel size 3 to map the 32-dim PPGs to 128-dim latent. We use the 2-layer Conformer [gulati2020conformer] before the upsampling layer with $4$ attention heads, 512 feed-forward dimension, and a convolution kernel size of 9. After the upsampling, we use a 2-layer Transformer with 4 attention heads and 512 feed-forward dimensions. We find this configuration yields the best outcome.
Our flow-matching decoder follows the same configuration as in [matchatts]. The external speaker embedding is of 256 dimensions. We use an embedding table of dimension 16 to represent the quantized pitch and concatenate the log periodicity with the pitch embedding as the soft voiced/unvoiced (V/UV) flag. External speaker embedding and pitch embedding with the V/UV flag are concatenated together as the conditional input to the decoder. During training, there is a $10\%$ probability for each batch that the PPG latent and the conditional input will be dropped to train the unconditioned model. During inference, we set the guidance strength $w=3$, and the sway coefficient is set to 1-1.
We train both the Matcha-TTS and our PPG2Speech model from scratch using the training set and evaluate on the test set and the unseen test set. Both models are trained with a vanilla flow-matching decoder in Matcha-TTS and a flow-matching decoder strengthened by CFG, to show the CFG's effect on the model's performance. All models are trained with a batch size of 32 for 500k updates on 1 Nvidia A100 GPU. The maximum learning rate is 1e-4, and we use a cosine annealing learning rate scheduler with 150k steps linear warmup to adjust the learning rate during training.
@@ 4.3.1 Objective Evaluation
The results of the objective evaluation are shown in Table 3. We use Speaker Encoder Cosine Similarity (SECS) to measure the similarity between the groundtruth speaker embedding and the speaker embedding extracted from the synthesized speech using the same SimAMResNet34 model in Section 4.1. We use a Wav2Vec2-large model [^2] https://huggingface.co/GetmanY1/wav2vec2-large-fi-150k-finetuned that is pretrained on 158k hours of unlabeled Finnish speech and fine-tuned on 4600 hours of Finnish speech to evaluate the intelligibility of the synthesized speech and use Character Error Rate (CER) as the objective metric. Furthermore, we use Mel-Cepstral Distortion (MCD) [chen2022v2c] and pitch mean absolute error (Pitch MAE) to evaluate the model's capability of handling unseen speakers and how well the model follows the given pitch, respectively.
@@
    CFG improves the model's generalization to unseen speakers.
@@
    Estimate the PPG using the same HMM-TDNN-medium model as Section 4.1 and use Montreal-Forced-Aligner [mcauliffe17_interspeech] to find the corresponding edited region. Calculate the PAC in Section 3.2 between the edited PPG and the estimated PPG in the edited region.